# How Memory Management Impacts LLM Agents: An Empirical Study of Experience-Following Behavior

### v1

@@ A.4 Coarse evaluator prompts
System Prompt You are a helpful AI assistant. Solve tasks using your coding and language skills. In the following cases, suggest python code (in a python coding block) or shell script (in a sh coding block) for the user to execute. 1. When you need to collect info, use the code to output the info you need, for example, browse or search the web, download/read a file, print the content of a webpage or a file, get the current date/time. After sufficient info is printed and the task is ready to be solved based on your language skill, you can solve the task by yourself. 2. When you need to perform some task with code, use the code to perform the task and output the result. Finish the task smartly. Solve the task step by step if you need to. If a plan is not provided, explain your plan first. Be clear which step uses code, and which step uses your language skill. When using code, you must indicate the script type in the code block. The user cannot provide any other feedback or perform any other action beyond executing the code you suggest. The user can’t modify your code. So do not suggest incomplete code which requires users to modify. Don’t use a code block if it’s not intended to be executed by the user. If you want the user to save the code in a file before executing it, put # filename: $<$filename$>$ inside the code block as the first line. Don’t include multiple code blocks in one response. Do not ask users to copy and paste the result. Instead, use ’print’ function for the output when relevant. Check the execution result returned by the user. If the result indicates there is an error, fix the error and output the code again. Suggest the full code instead of partial code or code changes. If the error can’t be fixed or if the task is not solved even after the code is executed successfully, analyze the problem, revisit your assumption, collect additional info you need, and think of a different approach to try. When you find an answer, verify the answer carefully. Include verifiable evidence in your response if possible. Reply "TERMINATE" in the end when everything is done.
@@
System Prompt You are a highly knowledgeable and rigorous judge for autonomous driving. You are judging a *short-horizon* trajectory (e.g., 6 steps). We only require the following: 1) The predicted trajectory should *generally* move towards or align with the goal. 2) It should stay within a drivable area (i.e., allowed region). 3) It should avoid collisions with other objects. Your output format: - First line: strictly output ’yes’ or ’no’. - Following lines: provide your reasoning (Chain-of-Thought is allowed). Be mindful that small lateral or partial forward movements can be acceptable as long as the overall direction is consistent with the planning target and safety requirements. Our coordinate system is such that the x-axis is lateral, and the y-axis is forward. Therefore, moving forward means an increase in y values. Be mindful that minor lateral adjustments or minimal forward movements are acceptable. If the y coordinate is increasing from step to step (and there’s no collision or out-of-lane), that may be considered a success.
User Prompt Below are the relevant information for this autonomous driving task: 1) Current state of the ego vehicle: {ego_prompts} 2) Perception of the environment: {perception_prompts} 3) Commonsense: {commonsense_mem} 4) Planning target: {planning_target} 5) Predicted trajectory: {pred_traj} Please decide if the predicted trajectory is successful under the above criteria, then provide your reasoning (you may use chain-of-thought). Remember: - First line of your answer: ’yes’ or ’no’ ONLY. - Following lines: your reasons or chain-of-thought.
@@
User Prompt You are an expert in identifying IoT attack categories. I will provide you with a problem and an answer that needs evaluation. Your task is to determine whether the given answer is correct or not. ### Problem: Based on the following features, determine the most likely attack type from the list below: ### Features: {problem} ### Answer to be Evaluated: {model_answer} - Respond with your judgement and explanation as following format. - First line: Respond with ’correct’ or ’incorrect’ only. - Following lines: Provide your reasoning or chain-of-thought. Your judgement:
@@ Appendix B Additional results
We present additional results in this section.
@@
Figure 10: Cumulative average input similarity vs. performance of different addition strategies over execution.
Figure 11: Addition results for misaligned memory reply on CIC-IoT Agent.
Figure 12: Addition results for misaligned memory reply on AgentDriver.
Appendix C Additional ablation results of AgentDriver
C.1 Experiments on different LLM backbones
@@
C.2 Error-free variant of history-based deletion
Figure 16: Comparison between history-based deletion and its error-free variants.
In Figure 16, we follow the procedure in Section 4.4 to plot the history-based deletion and its error-free variant. Surprisingly, we observe that around task index 1000, the combined deletion with strict addition surpasses the performance of its error-free variant. This suggests that history-based deletion can retain memory entries with outputs suitable for later reuse, corroborating the necessity of adopting this type of deletion.
C.3 Performance trend with noisy initialization
In this experiment, as shown in Figure 17, we initialize the memory bank in the same manner as the fixed memory bank setting described in the main experiment. Additionally, noise is introduced to the executions of the initial memories, following a normal distribution with a mean of 0 and a variance of 5.0. The experiment is conducted on a subset comprising the first 500 test samples out of the 2000 test samples used in the main experimental setup.
Figure 17: Average performance trend over time with initial memory that has noise with L2=5 on AgentDriver
C.4 Performance with different fixed memory sizes
We provide additional results on memory resource constraints with different memory limitation sizes.
Figure 18: Different limited sizes versus the performance of AgentDriver. The horizontal dashed lines are their corresponding unlimited variant performance.

### v2

@@ A.4 Coarse evaluator prompts
System Prompt You are an expert judge for Electronic Health Records (EHR) database queries and analysis. Your task is to evaluate whether the provided code solution and execution result are reasonable and correct for the given medical database question. You should assess the solution based on: 1) **Code Quality**: Does the code use appropriate database functions and follow logical steps? Be lenient about minor inefficiencies or alternative valid approaches. 2) **Result Reasonableness**: Does the execution result appear plausible and well-formatted for the medical context? 3) **Completeness**: Does the solution address the core question asked? Accept solutions that may miss minor constraints but solve the main problem. **Evaluation Guidelines:** - Focus on whether the solution would produce a reasonable answer in practice - Accept alternative valid approaches even if not optimal - Be lenient about missing minor date filters if the core logic is sound - Consider the medical context - some variation in results is normal - Only reject solutions that are fundamentally flawed or completely irrelevant **Available Tool Functions:** (1) Calculate(FORMULA), which calculates the FORMULA and returns the result. (2) LoadDB(DBNAME) which loads the database DBNAME and returns the database. The DBNAME can be one of the following: admissions, chartevents, cost, d_icd_diagnoses, d_icd_procedures, d_items, d_labitems, diagnoses_icd, icustays, inputevents_cv, labevents, microbiologyevents, outputevents,patients, prescriptions, procedures_icd, transfers. (Continued on next page…)
System Prompt (Continuation from previous page…)
(3) FilterDB(DATABASE, CONDITIONS), which filters the DATABASE according to the CONDITIONS. The CONDITIONS is a string composed of multiple conditions, each of which consists of the column_name, the relation and the value (e.g., COST<10). The CONDITIONS is one single string (e.g., "admissions, SUBJECT_ID=24971"). (4) GetValue(DATABASE, ARGUMENT), which returns the values of the column in the DATABASE. When there is no additional operations on the values, the ARGUMENT is the column_name in demand. If the values need to be returned with certain operations, the ARGUMENT is composed of the column_name and the operation (like COST, sum). Please do not contain " or ’ in the argument. (5) SQLInterpreter(SQL), which interprets the query SQL and returns the result. (6) Calendar(DURATION), which returns the date after the duration of time.
**Available Tables:** admissions, chartevents, cost, d_icd_diagnoses, d_icd_procedures, d_items, d_labitems, diagnoses_icd, icustays, inputevents_cv, labevents, microbiologyevents, outputevents, patients, prescriptions, procedures_icd, transfers **Knowledge of the Tables:** (1) Tables are linked by identifiers which usually have the suffix ’ID’. For example, SUBJECT_ID refers to a unique patient, HADM_ID refers to a unique admission to the hospital, and ICUSTAY_ID refers to a unique admission to an intensive care unit. (2) Charted events such as notes, laboratory tests, and fluid balance are stored in a series of ’events’ tables. For example the outputevents table contains all measurements related to output for a given patient, while the labevents table contains laboratory test results for a patient. (Continued on next page…)
System Prompt (Continuation from previous page…)
(3) Tables prefixed with ’d_’ are dictionary tables and provide definitions for identifiers. For example, every row of chartevents is associated with a single ITEMID which represents the concept measured, but it does not contain the actual name of the measurement. By joining chartevents and d_items on ITEMID, it is possible to identify the concept represented by a given ITEMID.
(4) For the databases, four of them are used to define and track patient stays: admissions, patients, icustays, and transfers. Another four tables are dictionaries for cross-referencing codes against their respective definitions: d_icd_diagnoses, d_icd_procedures, d_items, and d_labitems. The remaining tables, including chartevents, cost, inputevents_cv, labevents, microbiologyevents, outputevents, prescriptions, procedures_icd, contain data associated with patient care, such as physiological measurements, caregiver observations, and billing information. For different tables, they contain the following information: (1) admissions: ROW_ID, SUBJECT_ID, HADM_ID, ADMITTIME, DISCHTIME, ADMISSION_TYPE, ADMISSION_LOCATION, DISCHARGE_LOCATION, INSURANCE, LANGUAGE, MARITAL_STATUS, ETHNICITY, AGE (2) chartevents: ROW_ID, SUBJECT_ID, HADM_ID, ICUSTAY_ID, ITEMID, CHARTTIME, VALUENUM, VALUEUOM (3) cost: ROW_ID, SUBJECT_ID, HADM_ID, EVENT_TYPE, EVENT_ID, CHARGETIME, COST (4) d_icd_diagnoses: ROW_ID, ICD9_CODE, SHORT_TITLE, LONG_TITLE (5) d_icd_procedures: ROW_ID, ICD9_CODE, SHORT_TITLE, LONG_TITLE (6) d_items: ROW_ID, ITEMID, LABEL, LINKSTO (7) d_labitems: ROW_ID, ITEMID, LABEL (8) dianoses_icd: ROW_ID, SUBJECT_ID, HADM_ID, ICD9_CODE, CHARTTIME (Continued on next page…)
System Prompt (Continuation from previous page…)
(9) icustays: ROW_ID, SUBJECT_ID, HADM_ID, ICUSTAY_ID, FIRST_CAREUNIT, LAST_CAREUNIT, FIRST_WARDID, LAST_WARDID, INTIME, OUTTIME (10) inputevents_cv: ROW_ID, SUBJECT_ID, HADM_ID, ICUSTAY_ID, CHARTTIME, ITEMID, AMOUNT (11) labevents: ROW_ID, SUBJECT_ID, HADM_ID, ITEMID, CHARTTIME, VALUENUM, VALUEUOM
(12) microbiologyevents: RROW_ID, SUBJECT_ID, HADM_ID, CHARTTIME, SPEC_TYPE_DESC, ORG_NAME (13) outputevents: ROW_ID, SUBJECT_ID, HADM_ID, ICUSTAY_ID, CHARTTIME, ITEMID, VALUE (14) patients: ROW_ID, SUBJECT_ID, GENDER, DOB, DOD (15) prescriptions: ROW_ID, SUBJECT_ID, HADM_ID, STARTDATE, ENDDATE, DRUG, DOSE_VAL_RX, DOSE_UNIT_RX, ROUTE (16) procedures_icd: ROW_ID, SUBJECT_ID, HADM_ID, ICD9_CODE, CHARTTIME (17) transfers: ROW_ID, SUBJECT_ID, HADM_ID, ICUSTAY_ID, EVENTTYPE, CAREUNIT, WARDID, INTIME, OUTTIME **Correct Examples:** {EHRAgent_4Shots_Knowledge} Now, please judge whether the provided solution is reasonable and correct. Please assume that the solution is bug-free and does not contain any syntax errors. ### Execution to be judged: Question:{question} Knowledge:{knowledge} Solution:{code} ### Execution Result: {execution_result} ### Your Task: Evaluate whether this solution is reasonable and correct. Consider the medical context, database logic, and result plausibility. **Output Format:** - First line of your answer: ’yes’ or ’no’ ONLY. - Following lines: briefly provide your reasoning and analysis. Your evaluation:
@@
System Prompt You are a highly knowledgeable and rigorous judge for autonomous driving. You are judging a *short-horizon* trajectory (e.g., 6 steps). We only require the following: 1) The predicted trajectory should *generally* move towards or align with the goal. 2) It should stay within a drivable area (i.e., allowed region). 3) It should avoid collisions with other objects. Your output format: - First line: strictly output ’yes’ or ’no’. - Following lines: provide your reasoning (Chain-of-Thought is allowed). Be mindful that small lateral or partial forward movements can be acceptable as long as the overall direction is consistent with the planning target and safety requirements. Our coordinate system is such that the x-axis is lateral, and the y-axis is forward. Therefore, moving forward means an increase in y values. Be mindful that minor lateral adjustments or minimal forward movements are acceptable. If the y coordinate is increasing from step to step (and there’s no collision or out-of-lane), that may be considered a success. Below are the relevant information for this autonomous driving task: 1) Current state of the ego vehicle: {ego_prompts} 2) Perception of the environment: {perception_prompts} 3) Commonsense: {commonsense_mem} 4) Planning target: {planning_target} 5) Predicted trajectory:
User Prompt {pred_traj} Please decide if the predicted trajectory is successful under the above criteria, then provide your reasoning (you may use chain-of-thought). Remember: - First line of your answer: ’yes’ or ’no’ ONLY. - Following lines: your reasons or chain-of-thought.
@@
User Prompt You are a strict evaluator for IoT traffic classification answers. You should infer the most likely label from the provided flow-level features, then judge whether the model’s answer is CORRECT or INCORRECT. ### Judging Criteria The Model Answer is CORRECT if its predicted label exactly matches your inferred gold label, else INCORRECT. ### Judging Steps 1) Carefully analyze key fields and reason your inferred gold label. 2) Compare the Model Answer’s label to your gold label. ### Problem: Based on the following features, determine the most likely traffic type from the list below: ### Features: problem ### Model Answer: {generated_answer} - Respond with your judgement and explaination as following format. - First line: Respond with ’CORRECT’ or ’INCORRECT’ only. - Following lines: Provide your reasoning or chain-of-thought. Your judgement:
A.5 Additional setup for task distribution shift
For task distribution shift, we obtain the embedding vectors for all test queries using OpenAI text-embedding-3-large encoder for EHRAgent, and the original input vector computation method from AgentDriver. We then apply a Gaussian Mixture Model (GMM) to cluster the embeddings into three groups – each group will likely have a distinguished task distribution. Finally, we reorder the test queries based on their assigned cluster labels to process the three task groups sequentially. This procedure forces a pronounced shift in the task distribution.
@@ Appendix B Additional results
B.1 Additional results on addition experiments
We present additional results for the memory addition experiments in this section.
@@ B.1 Additional results on addition experiments
B.2 Experiments on different LLM backbones
@@
B.3 Error-free variant of history-based deletion
Figure 13: Comparison between history-based deletion and its error-free variants. Coarse here represents C1 evaluator.
In Figure 13, we follow the procedure in Section 3.4 to plot the history-based deletion and its error-free variant. Surprisingly, for AgentDriver, we observe that around task index 1000, the combined deletion with strict addition surpasses the performance of its error-free variant. This suggests that history-based deletion can retain memory entries with outputs suitable for later reuse, corroborating the necessity of adopting this type of deletion.
B.4 History-based deletion on different evaluators
Figure 14: Performance comparison after applying history-based deletion with different evaluators.
In Figure 14, we present performance changes after history-based deletion on RegAgent and CIC-IoT Agent with different evaluators.
B.5 Comparison between deleted records and retained records
Figure 15: Comparison between deleted and retained records for RegAgent under history-based deletion. The top figure shows results with Coarse 1.2, and the bottom figure shows results with Coarse 1.4. Lower absolute error represents better execution quality.
Figure 16: Comparison between deleted records and retained records for AgentDriver. Lower error represents better execution quality.
Table 4: Correctness rate (%) of deleted records and retained records for EhrAgent and CIC-IoT Agent with coarse evaluation.
Agent Category 4o-mini 4.1-mini 4.1-mini-FT EhrAgent Retained 44.1 49.1 54.8 Deleted 36.3 32.1 48.2 CIC-IoT Retained 78.9 72.2 86.6 Deleted 56.7 55.1 61.0
In this section, we analyze the intrinsic quality differences between retained and deleted memory records. Figure 15 presents results for the Coarse 1 and Coarse 2 evaluators under history-based deletion, while Figure 16 shows results across all evaluators for AgentDriver. Interestingly, when using GPT-4o-mini as the evaluator, the retained memory in AgentDriver exhibits lower average quality than the deleted records. We further report the correctness rates with respect to ground truth for EhrAgent and CIC-IoT Agent, as shown in Table 4.
B.6 Results under memory resource constraints.
We provide the results of the agent running under memory resource constraints in Figure 17. We also provide additional results on memory resource constraints with different memory limitation sizes in Figure 18.
Figure 17: Comparison of unlimited memory size versus a limited memory size for strict and coarse selective addition with combined deletion. Coarse here denotes the C1 evaluator.
Figure 18: Different limited sizes versus the performance of AgentDriver. The horizontal dashed lines are their corresponding unlimited variant performance. Coarse here denotes the C1 evaluator.